OpenStudy (anonymous):

Write the equation of the line that passes through (–2, 1) and is perpendicular to the line 3x – 2y = 5. HELP!

OpenStudy (anonymous):

first put the original equation in the form of y =mx+b m being the slope, and b being the y-intercept

OpenStudy (anonymous):

to find perpendicular, the slope would be inverse with one another with different signs

OpenStudy (anonymous):

I am not good with slope, that's why I'm asking for help.

OpenStudy (anonymous):

I also have 2 more questions, if you could help me with those after this as well?

OpenStudy (anonymous):

the slope is already given, you just need to put 3x - 2y =5 in to y = mx+b form, and your slope would be the m in the equation

OpenStudy (anonymous):

I don't know how to do that, that's what I'm saying.

OpenStudy (anonymous):

 now you have it in the form of y=mx+b